OpenStudy (kaylee):

What is the solution to the equation 2(2-j)= 4(j+10) ? j =

OpenStudy (revlogic):

2(2 - j)= 4(j+10) => 4 -2j = 4j + 40 => -6j = 36 => j = 36/-6 => -6

OpenStudy (kaylee):

Okay thank you

OpenStudy (revlogic):

you're welcome :)

OpenStudy (alivejeremy):

Answer: j=−6

OpenStudy (alivejeremy):

First, simplify both sides of the equation.

OpenStudy (alivejeremy):

Next, subtract 4 from both sides.

OpenStudy (alivejeremy):

Then and Last divide both sides by -6.

OpenStudy (kaylee):

Thank you